# Dynamical model for black hole to white hole transitions

### v1

@@ I Introduction
A complete theory of quantum gravity remains elusive. However, there is broad agreement on some of its expected features. One such expectation, perhaps the most important, is singularity resolution and the final state of black holes. There are three broad approaches to addressing this question: propose phenomenological metrics using general arguments [1, 2, 3, 4, 5, 6] and study their properties without recourse to any specific theory; quantize the Schwarzschild black hole [7, 8, 9, 10, 11, 12], and seek static or dynamical solutions of proposed effective equations with matter that contain quantum gravity corrections [13, 14, 15, 16, 17, 18, 19, 20].
@@ II Black hole to White hole transition metric
$m(r,t)=M_{0}\left[1+\tanh\left(\frac{r-r_{0}-v(r,t)t}{\alpha\ l_{0}}\right)\right]^{a}\tanh\left[\left(\frac{r}{l_{0}}\right)\right]^{b}.$ (3)
@@ II.1 Transient horizons
The formation and evolution of horizons is demonstrated by computing $\theta_{\pm}(r,t)$. The roots of these functions give the locations of outer and inner horizons as a function of $t$. Fig. 1 shows typical plots of $r\theta_{\pm}/2$ for the parameter values $M_{0}=4,\ a=2,\ b=3,\ r_{0}=5,\ t_{0}=0$, with all other parameters set to unity. There are no roots in the first frame $(t=-1200)$ indicating no horizons; the next two frames show roots of $\theta_{+}$ indicating the formation of a black hole with the outer horizon fixed at $r=2M_{\text{ADM}}=32$ (for the chosen parameters); the inner horizon moves inward with the matter flow; the fourth frame ($t=0$) is the transition point with no roots and therefore no horizons; the next frame ($t=200$) shows an inner and outer white hole horizon at $r=32$, followed by disappearance of all horizons in the last frame ($t=1200$).
Figure 1: Null expansion scalars $\theta_{\pm}$ that exhibit a transition from a black hole ($\theta_{+}<0$ for some range of $r$ and $\theta_{-}<0$ for all $r$) to a white hole ($\theta_{-}>0$ for some range of $r$ and $\theta_{+}>0$ for all $r$) at the times indicated: a BH of radius $r=32$ forms at $t=-800$ and disappears at $t=0$; a WH of radius $r=32$ has formed at $t=20$ and disappears at $t=1200$.
@@
Another view of this feature is provided in Fig. 2 which shows the BH and WH regions in the $r-t$ plane: the left frame shows the inner (blue) and outer (red) horizons, and the peak value of the matter density (6)(green); the right frame is a close-up version that resolves the separation between the BH and WH regions near the matter bounce time $t=0$; also shown are a sampling of light cones. These figures show several notable features: (i) the outer BH horizon turns spacelike, and moves radially inward to meet the inner BH horizon near $t\approx 0.5$; (ii) the matter density leaves the BH region as the inner and outer BH horizons annihilate near $t=-0.5$, and then enters the WH region while moving in a timelike manner; (iii) during the relatively small time interval between $t=-0.5$ and $t=0.5$, there are no horizons, and one may interpret the metric as that of a transient remnant emerging from the black hole region and then transitioning into a white hole.
@@ III Summary and discussion
We proposed and studied a smooth dynamical spherically symmetric non-singular asymptotically flat metric that describes a BH $\rightarrow$ WH transition. The metric may be viewed as a generalization of [21] where there is matter bounce and black hole evaporation, but no white hole formation. In comparison to other works such as [22], the metric we give is constructed by “following the matter” using a “designer” mass function in PG coordinates, rather than by gluing together several metrics to achieve the desired outcome. Although the metric captures several desirable features that might emerge from a theory of quantum gravity, it remains to be seen whether it can arise as a solution in an effective theory with quantum gravity corrections.
There are a few specific features of the geometry that are noteworthy: (i) the BH and WH horizons do not intersect at a point, unlike what is proposed in models that use junctions conditions; (ii) the matter and inner horizons move together and are timelike in some regions and spacelike in others; (iii) matter is superluminal where the BH region ends and the WH region begins (Fig. 2); (iv) the DEC is violated in these regions.

### v2

@@ I Introduction
A complete theory of quantum gravity remains elusive. However, there is broad agreement on some of its expected features. One such expectation, perhaps the most important, is singularity resolution and the final state of black holes. There are three broad approaches to addressing this question: propose phenomenological metrics using general arguments [19, 2, 17, 11, 5, 6, 28, 9] and study their properties without recourse to any specific theory; quantize the Schwarzschild black hole [26, 29, 8, 3, 13, 16], and seek static or dynamical solutions of proposed effective equations with matter that contain quantum gravity corrections [24, 33, 25, 32, 22, 21, 14, 15].
@@ II Black hole to White hole transition metric
$m(r,t)=M_{0}\left[1+\tanh\left(\frac{r-r_{0}-v(r,t)\,(t-t_{0})}{\alpha\ l_{0}}\right)\right]^{a}\tanh\left[\left(\frac{r}{l_{0}}\right)\right]^{b}.$ (3)
@@ II.1 Transient horizons
The formation and evolution of horizons is demonstrated by computing $\theta_{\pm}(r,t)$. The roots of these functions give the locations of outer and inner horizons as a function of $t$. Fig. 1 shows typical plots of $r\theta_{\pm}/2$ for the parameter values $M_{0}=4,\ a=2,\ b=3,\ r_{0}=5,\ t_{0}=0$, with all other parameters set to unity. There are no roots in the first frame (t=1200)(t=-1200) indicating no horizons; the next two frames show roots of $\theta_{+}$ indicating the formation of a black hole with the outer horizon fixed at $r=2M_{\text{ADM}}=32$ (for the chosen parameters); the inner horizon moves inward with the matter flow; the fourth frame ($t=0$) is the transition point with no roots and therefore no horizons; the next frame ($t=200$) shows an inner and outer white hole horizon at $r=32$, followed by disappearance of all horizons in the last frame ($t=1200$).
Figure 1: Null expansion scalars $\theta_{\pm}$ that exhibit a transition from a black hole ($\theta_{+}<0$ for some range of $r$ and $\theta_{-}<0$ for all $r$) to a white hole ($\theta_{-}>0$ for some range of $r$ and $\theta_{+}>0$ for all $r$) at the times indicated: a BH of radius $r=32$ forms at t=800t=-800 and disappears at $t=0$; a WH of radius $r=32$ has formed at $t=20$ and disappears at $t=1200$.
@@
Another view of this feature is provided in Fig. 2 which shows the BH and WH regions in the rtr-t plane: the left frame shows the inner (blue) and outer (red) horizons, and the peak value of the matter density (6)(green); the right frame is a close-up version that resolves the separation between the BH and WH regions near the matter bounce time $t=0$; also shown are a sampling of light cones. These figures show several notable features: (i) the outer BH horizon turns spacelike, and moves radially inward to meet the inner BH horizon near $t\approx 0.5$; (ii) the matter density leaves the BH region as the inner and outer BH horizons annihilate near t=0.5t=-0.5, and then enters the WH region while moving in a timelike manner; (iii) during the relatively small time interval between t=0.5t=-0.5 and $t=0.5$, there are no horizons, and one may interpret the metric as that of a transient remnant emerging from the black hole region and then transitioning into a white hole.
@@ III Summary and discussion
We proposed and studied a smooth dynamical spherically symmetric non-singular asymptotically flat metric that describes a BH $\rightarrow$ WH transition. The metric may be viewed as a generalization of [20] where there is matter bounce and black hole evaporation, but without white hole formation. In comparison to other works such as [18], the metric we give is constructed by “following the matter” using a “designer” mass function in PG coordinates, rather than by gluing together several metrics to achieve the desired outcome. Although the metric captures several desirable features that might emerge from a theory of quantum gravity, it remains to be seen whether it can arise as a solution in an effective theory with quantum gravity corrections. This requires deriving effective equations for gravitational collapse in spherical symmetry. As noted above, such equations have been derived for pressureless dust in LQG, but their solutions do not predict a BH to WH transition. The question is whether collapse of matter with pressure, such as a perfect fluid or scalar field gives rise to WH formation after a bounce. In this respect, the metric we have presented can be viewed as a phenomenological model that describes a plausible dynamics that could arise from an actual effective theory. As noted above, this is similar in spirit to several other mostly static metrics for wormholes and dynamical metrics using junction conditions.
@@
There are a few specific features of the geometry that are noteworthy: (i) the BH and WH horizons do not intersect at a point, unlike what is proposed in models that use junctions conditions; (ii) the matter and inner horizons move together and are timelike in some regions and spacelike in others; (iii) matter is superluminal within the BH and WH, but becomes timelike in the gap region between them (Fig. 2); (iv) the DEC is violated only in the spacetime region where the bounce occurs, as expected in any bounce scenario, but the energy density $\rho$ remains positive everywhere; in particular, the DEC is not violated at sufficiently large radial coordinate.
On the first of these observations, it is useful to contrast with the “conventional” conjectured picture of a transition where the BH and WH horizons intersect at a point. In contrast, in metric we present, there is a short-lived gap between the BH and WH regions. We think this is a required feature of a dynamical transition because the condition for a single point of contact $(t_{c},r_{c})$ between the BH ($\theta_{-}<0,\ \theta_{+}<0$) and WH ($\theta_{-}>0,\ \theta_{+}>0$) regions would require $\theta_{+}(t_{c},r_{c})=\theta_{-}(t_{c},r_{c})=0$, which from (7) leads to an obvious contradiction. We note further that the qualitative features of the BH to WH transition are unaffected and robust if other choices of smooth sign switching of the shift function are used.
It is possible that there are astrophysical implications of the metrics we proposed. A first step in this direction is to study geodesics to see if there are possible observational signatures of the transition, especially of the short-lived “remnant” between the BH and WH regions. That remnants might provide a contribution to dark matter, along with primordial black holes, has been discussed in the literature [10, 4].